Suppose you mix 25.0 mL 0.50 M Sr(OH)2 with 40.0 mL of 0.50 M HCl. Calculate the pH of the resulting solution.
Answer #1

[OH-] = 2*[Sr(OH)2]

= 2*0.50 M

= 1.0 M

[H+] = [HCl]

= 0.50 M

Given:

M(H+) = 0.5 M

V(H+) = 40 mL

M(OH-) = 1 M

V(OH-) = 25 mL

mol(H+) = M(H+) * V(H+)

mol(H+) = 0.5 M * 40 mL = 20 mmol

mol(OH-) = M(OH-) * V(OH-)

mol(OH-) = 1 M * 25 mL = 25 mmol

We have:

mol(H+) = 20 mmol

mol(OH-) = 25 mmol

20 mmol of both will react

remaining mol of OH- = 5 mmol

Total volume = 65.0 mL

[OH-]= mol of base remaining / volume

[OH-] = 5 mmol/65.0 mL

= 7.692*10^-2 M

use:

pOH = -log [OH-]

= -log (7.692*10^-2)

= 1.1139

use:

PH = 14 - pOH

= 14 - 1.1139

= 12.8861

Answer: 12.89